OpenStudy (anonymous):

A copper wire (density = 8.96 g/cm^3) has a diameter of 0.35 mm. If a sample of this copper wire has a mass of 10. g, how long is the wire?

OpenStudy (anonymous):

Take HL math.

OpenStudy (anonymous):

Okay so you are given density, mass, and the diameter of a copper wire that means that you can relate the information given to volume of the wire and utilize density formula.

OpenStudy (anonymous):

D = m/v, Volume = mass/density

OpenStudy (anonymous):

Now think of the wire as a cylinder so (pi)(r^2)L

OpenStudy (anonymous):

So (pi)(r^2)L = 10g/(8.96g/cm^3)

OpenStudy (anonymous):

pi = 3.14....(more precise using this function on your calculator) r = (diameter)/2 = 0.35/2 = 0.175

OpenStudy (anonymous):

So now the only unknown left in your equation is L which represents the length of the wire (or how long the wire is)

OpenStudy (anonymous):

Before we proceed you have make sure that your units of length are the same so you can convert the radius 0.175 mm to cm

OpenStudy (anonymous):

radius = 0.0175 cm

OpenStudy (anonymous):

L = 10g/(8.96g/cm^3)/(0.0175cm)^2/(3.14) L = 1160.02145cm Do you follow?

OpenStudy (anonymous):

thank you so much! very clear

OpenStudy (anonymous):

great! you're welcome